# Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities

### v2

@@ III Disorder Operator
Once the transition points of each $\Delta$ are determined, we now carry out the analysis of the disorder operator upon them. As a non-local observable, disorder operator is defined as the expectation value of a symmetry transformation applied to a finite region in the statistical or quantum many-body systems of interest Wegner (1971); Kadanoff and Ceva (1971); Fradkin (2017); Nussinov and Ortiz (2009a, b). The design and implementation of the disorder operator and the analysis of its finite size scaling behavior have been successfully carried out in the situations of spontaneous symmetry breaking phase, quantum critical points, the symmetric phases with topological orders, symmetric mass generation transition and even the free fermion surface and interacting quantum critical Fermi surface systems Zhao et al. (2021); Wu et al. (2021a); Wang et al. (2021); Wu et al. (2021b); Chen et al. (2022); Wang et al. (2022); Jiang et al. (2023); Liu et al. (2023, 2024); Cai and Cheng (2024); Wu (2024).
@@
We have further computed a related quantity – the bipartite spin fluctuation Song et al. (2011). At $\theta\to 0$ limit, the scaling of the disorder operator $-\ln|X_{M}|$ has a similar behavior to that of the bipartite spin fluctuations, which is defined as
@@
Figure 4: Second Rényi entropy $S^{(2)}_{A}$ for $L\times L$ lattice with subsystem size of $L/2\times L$ deep in the AFXY phase of EPJQ model at $\Delta=0$ and $q=0.3$. (a) $S^{(2)}_{A}$ versus system size $L$. The black curve is a fit to $S^{(2)}_{A}=aL+b\ln(L)+c$ and the fitting result is $S^{(2)}_{A}=0.1338(8)L+0.497(11)\ln(L)+0.0046(166)$. (b) $S^{(2)}_{A}-aL$ vs. $\ln(L)$ for different system sizes where we choose the fitted value $a=0.1338$. The black curve is a straight line of $y=0.497x+0.0046$.
@@
Figure 6: Fitted log-coefficient $b$ according to Eq. (10) with finite-size data of $\rho_{s}(L)$ and $I(L)$ obtained from QMC simulations at the phase transition points of EPJQ model for different $\Delta$. By considering the finite size values of $\rho_{s}(L)$ and $I(L)$, the obtained $b$ indeed gets enhanced, although with larger error bars. The black dotted lines are guides for eyes for the log coefficient of three Goldstone modes from the SO(4) symmetry-breaking (for $\Delta<1$) and four Goldstone modes from the SO(5) symmetry-breaking (for $\Delta=1$) cases.
@@ IV Entanglement Entropy
Next we investigate the EE at the transition points of various $\Delta$ of Eq. (1). To this end, we employ the nonequilibrium increment method Alba (2017); D’Emidio (2020); Zhao et al. (2022a, b) within the framework of SSE-QMC simulation Sandvik (1999); Syljuåsen and Sandvik (2002) to determine the second Rényi entropy of the EPJQ model at phase transition points for various values of $\Delta$. Rényi entanglement entropy is defined as $S_{A}^{(n)}=\frac{1}{1-n}\ln\operatorname{Tr}\rho_{A}^{n}$ which can be reexpressed in the form of $S_{A}^{(n)}=\frac{1}{1-n}\ln\frac{Z_{A}^{(n)}}{Z^{(n)}}$ according to the replica trick Calabrese and Cardy (2004). The nonequilibrium method is based on Jarzynski’s equality Jarzynski (1997), which relates the free energy difference between two systems with the total work done during a tuning process from one system to another. We regard the partition functions $Z_{A}^{(n)}$ and $Z^{(n)}$ as those of two different physical systems, then it is natural to apply the Jarzynski’s equality and design a tunning process between the two systems to calculate the Rényi entropies Alba (2017); D’Emidio (2020). In practice, we follow the incremental version Zhao et al. (2022a, b) of Ref. D’Emidio (2020), and it can overcome the obstacles that the EE is in general an exponential observable Zhang et al. (2024); Zhou et al. (2024). We conduct the simulation on a $L\times L$ square lattice with periodic boundary conditions, and fix with $\beta=L$ and choose the subregion A to be a $L\times L/2$ cylinder defined same as region $M$ in the right inset of Fig. 1 (a), which has no sharp corners on the entanglement boundary.
For $\Delta<1.0$, in the AFXY phase, the system spontaneously breaks the SO(2) spin symmetry and possesses one Goldstone mode. In this case, EE is expected to scale as Metlitski and Grover (2015)
@@
where $N_{g}$ represents the number of Goldstone modes. Eq. (9) has been extensively verified for both XY phase with one Goldstone mode and Néel ordered states with two Goldstone modes respectively Helmes and Wessel (2014); Kulchytskyy et al. (2015); D’Emidio (2020); Zhao et al. (2022b); Song et al. (2023a); Deng et al. (2023b); Zhou et al. (2024); D’Emidio et al. (2024) . We first test our algorithm in the AFXY phase of the EPJQ model at $\Delta=0$ and $q=0.3$ deep in the AFMXY phase, with the results displayed in Fig. 4. We measure the second order Rényi EE for the $L\times L/2$ subsystem $A$ on a $L\times L$ square lattice with smooth boundary. The system sizes utilized are $L=8,12,16,\ldots,32$ and the temperature is fixed at $\beta=L$. As shown in Fig. 4(b), by fitting the finite-size EE data to Eq. (9), we observe a good agreement between our data with the scaling form of Eq. (9). Moreover, the log-coefficient obtained from curve fitting is $\frac{N_{G}}{2}=0.497\pm 0.011$, aligning well with the expectation of existence of one Goldstone mode, which corresponds to spontaneous SO(2) symmetry breaking. The analysis of the fitting quality can be found in the Sec. II in SM sup .
@@
We then proceed to the phase transition points. Our extrapolated finite order parameter data in Fig. 1(b) shows that the first-order behavior is more pronounced as $\Delta$ decreases from 1 to 0. However, the relation between the scaling of entanglement entropy (EE) with the strengthen of first-order behavior remains unexplored. We measure the second Rényi EE at the transition points in Table. 1 for various values of $\Delta$ and system sizes $L=8,12,16,\ldots,44$ and fix $\beta=L$. The subsystem A is again chosen to be $L\times L/2$ cylinder withour corners on the boundary. As presented in Fig. 5, for all examined values of $\Delta$, we observe a clear logarithmic correction to the area law term in the scaling of EE (Eq. (9)). Additionally, as $\Delta$ decreases and the first-order behavior strengthens, the fitted log-coefficient $b$ increases. Interestingly, for $\Delta=0.1$ and $\Delta=0.3$, the fitted value of $b$ is even consistent with 0.5, which corresponds to the existence of a single Goldstone mode, the same as the scaling of EE deep inside the AFXY phase in Fig. 4. Our findings suggest that at least for $\Delta\leq 0.3$ the log-coefficient $b$ characterizes a evident first order transition with SO(2) symmetry breaking associated with one Goldstone mode.
At $\Delta=1.0$, where the system reduces to the isotropic JQ3 model, it has been discovered that even with a smooth entanglement boundary, the system exhibits anomalous positive logarithmic subleading corrections, resembling the Goldstone mode contribution, while the coefficient is only around 0.2-0.3 Song et al. (2023a, b). In addition, a recent study Deng et al. (2024) considers carefully the finite-size effects in scaling form for Goldstone mode phase, and apply the modified fitting function
$S^{(2)}_{A}(L)=aL+\frac{n_{G}}{2}\ln\Big{[}\big{(}\rho_{s}(L)L\big{)}^{1/2}\big{(}I(L)L\big{)}^{1/2}\Big{]}+c,$ (10)
where $\rho_{s}(L)$ is the finite-size spin stiffness and $I(L)$ the finite-size transverse spin susceptibility which are measured separately in QMC simulations and substituted in the above relation to fit the finite-size EE data. Through this procedure, the log-coefficient is fitted to be around 2 which corresponds to four Goldstone modes Deng et al. (2024). This is used to suggest that the system breaks the emergent SO(5) symmetry at the transition point, and this process can be well detected by the scaling of EE even at relatively small system sizes Deng et al. (2024).
We followed this procedure, calculated the finite-size spin stiffness $\rho_{s}(L)$ and transverse spin susceptibility $I(L)$, and substituted them in Eq. (10), same as in Ref. Deng et al. (2024) to fit our EE data. However, as shown in Fig. 6, we find it is difficult to observe the same convergence of log-coefficient to 2 at $\Delta=1.0$ or to 1.5 at $\Delta<1.0$, which would suggest the breaking of emergent SO(5) (four Goldstone modes) and SO(4) (three Goldstone modes) symmetries, respectively. In fact, at $\Delta=1.0$ the last two data points with larger $L_{\text{min}}$ are quite close to $b=2$, although the last data point slightly deviates below $b=2$.
For $\Delta<1.0$, except for $\Delta=0.3$, the last data points all seem to be consistent with $b=1.5$. However, within our limited data quality, we do not observe the stability of fitted results with respect to $L_{\text{min}}$ for any $\Delta$. It is possible that our results can be affected by stronger finite-size effects as we fixed $\beta=L$ to approach the thermodynamic limit, while Ref. Deng et al. (2024) used the projector quantum Monte Carlo Sandvik (2010, 2005) to simulate the ground state of the system and fix the projection power to be $m=L^{3}$. Another possible explanation is, as we observed at $\Delta<1.0$ the first order is enhanced, the emergent $SO(4)$ behavior can be weakened and broken into remaining VBS order and Goldstone phase at small system sizes. In this case, within the system sizes we considered, an evident first-order behavior with only one Goldstone mode might be mixed with the $SO(4)$ symmetry breaking behavior with three Goldstone modes. Then the fitted log coefficient with Eq. (10) for $\Delta<1.0$ could drift significantly from 1.5 to 0.5 as system sizes grow. Therefore our fitting results with Eq. (10) are consistent with Ref. Deng et al. (2024) for JQ3 model and show the capability of EE in characterizing the emergent symmetry and weakly first order behaviors at the complicated DQCPs.
@@ V Discussion
Our work shows the two non-local measurements – the disorder operator and entanglement entropy – consistently exhibit anomalous scaling behaviors at the easy-plane deconfined quantum criticalities. By adjusting the anisotropy from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$), we find that as the first-order nature of the transition is amplified, the anomalous behavior for both disorder operator and entanglement entropy becomes more apparent, resembling contributions from Goldstone modes. Interestingly, when $\Delta\leq 0.3$, the log coefficient in EE seem to converge to $b=0.5$ which is consistent with one Goldstone mode and SO(2) symmetry breaking at the first-order transition points. Additionally, when applying the finite-size fitting form of Goldstone mode phase (Eq. (10)) to fit our EE data, we obtain the enhanced log coefficients close to that of emergent SO(5) symmetry breaking at $\Delta=1$ and emergent SO(4) symmetry breaking at $\Delta<1$. However, the log coefficients do not show the good convergence behavior, as shown in Ref. Deng et al. (2024), which maybe explained by insufficient data quality or very weak emergent SO(4) symmetry at $\Delta<1$ due to the strengthening of first order behavior. Our work thus provides strong evidence that the observed anomalous scaling behaviors of the entanglement measurements (disorder operator and EE) indeed come from the weakly-first-order nature of DQCPs of JQ model realizations both at $\Delta=1$ and $\Delta<1$.
Note that our work and previous studies Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024); D’Emidio and Sandvik (2024); Deng et al. (2024) on the EE scaling of smooth cut on a square lattice are not contradicted with the recent work by D’Emidio and Sandvik D’Emidio and Sandvik (2024) where they consider a $\frac{\pi}{4}$-degrees-tilted square lattice and observe the normal scaling of EE obeying the requirements of CFTs. On the contrary, the cut dependence of EE might support the transition is indeed weakly-first order and the tilted cut and standard cut possibly just capture the two sides of the JQ-type DQCP story. On the one hand, it is a weakly first transition which can be detected by the standard cut, and on the other hand the system might be close to a real continuous transition so that the tilted cut captures some essential information of the CFT at small system sizes. In conclusion, scaling of these nonlocal observables have proven to be both powerful and sensitive tools in characterising this kind of weakly first-order transitions Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a); Deng et al. (2024), and it provides strong evidence that the JQ type DQCP is weakly first order however very close to a real continuous transition Takahashi et al. (2024), possibly of multicritical type related with the SO(5) model with Wess-Zumino-Witten term Zhao et al. (2020); Chen et al. (2023); Chester and Su (2024); Takahashi et al. (2024); Chen et al. (2024) or the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023).
Future research directions may include investigating the disorder operator and entanglement entropy scaling of JQn models with $n>3$, which have been shown to exhibit stronger first-order transitions Takahashi and Sandvik (2020); Takahashi et al. (2024) as $n$ increases, similar to EPJQ models. Another avenue of interest is to explore the entanglement entropy scaling of an absolutely strong first-order transition, which involves a mixture of two phases with classical probabilities. Moreover, multipartite entanglement Osborne and Nielsen (2002); Javanmard et al. (2018); Wang et al. (2024); Parez and Witczak-Krempa (2024) and higher-order Rényi entropies, entanglement negativity and entanglement spectrum Wu et al. (2020); Wang and Xu (2023); Yan and Meng (2023) with QMC simulations at the DQCP also present intriguing topics for future investigation.
@@ Acknowledgement
We thank Meng Cheng, Cenke Xu, Anders Sandvik, Senthil Todadri, Subir Sachdev for valuable discussions on the related topic. NVM and ZYM acknowledge the earlier insightful discussions on the transitions of EPJQ model with Arnab Sen and Anders Sandvik. JRZ and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037- 22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22), the GD-NSF (No. 2022A1515011007) and the HKU Seed Funding for Strategic Interdisciplinary Research. Y.C.W. acknowledges the support from Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003), and the support from the High-Performance Computing Centre of Hangzhou International Innovation Institute of Beihang University. NVM acknowledge the National Natural Science Foundation of China (No. 12004020) and the Fundamental Research Funds for the Central Universities. We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.
Supplemental Material for
"Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities"
I The criticality for different $\Delta$ in the EPJQ models
@@
Figure S1: (Upper panels) The crossing points $q_{c}(L)$ obtained from the curve of dimensionless Binder cumulants $U_{M}$(blue) and $U_{D}$(red) versus $q$ close to the phase transition points for two system sizes $L$ and $2L$. The red/blue curve is the fitting function with the scaling form of $q_{c}(L)$ in Eq. (5) excluding the crossings for small sizes with $\chi^{2}$ of the fitting becomes close to one. The data points presented at $01/L=$ 0 are the y-intecepts of fitting curves with the fitting errors which are given in Tab.1. (Lower panels) The value of order parameters for both AFXY and VBS phases at $q_{c}(L)$ obtained with corresponding Binder cumulants versus $1/L$ in the EPJQ models with different $\Delta$ for all computing sizes $L$. The red/blue curve is the second-order polynormial fitting function of all the data points with $\chi^{2}$ of the fitting close to one. The data points presented at $01/L=$ 0 are the y-intecepts of the fitting curves with errors illustrated in Fig. 1.
@@
The results in (b) of Fig. 1 in the main text present that the order parameters for both the AFXY and VBS phases converge to a non-zero value when $L$ goes to infinity for different $\Delta$ in the EPJQ models, which implies the possibility of first-order phase transitions in the EPJQ model. In this section we present the detail procedure in obtaining those converged order parameters at thermodynamic limit in Fig. 1(b). The binder cumulants $U_{M}$ and $U_{D}$ defined in Eq. (4) for the AFXY and VBS phases correspondingly are two dimensionless quantities that are commonly chosen in the FSS of calculating critical points and correlation length exponents. In this paper we study the critical behavior of the EPJQ model with different anisotropic value $\Delta$ with the help of $U_{M}$ and $U_{D}$ all the crossing points $q_{c}(L)$ of two simulated sizes $L$ and $2L$ got from $U_{M}$ and $U_{D}$ are shown in the first row in Fig. S1 for those different $\Delta$. Using the fitting form in Eq. (5) two $q_{c}(\infty)$ can be got, which are listed in Tab.1 in the main text. At each crossing point $q_{c}(L)$ we also calculate the square of order parameters defined in Eq. 2 and Eq. 3 as $M_{x}^{2}(L)$ and $D^{2}(L)$ illustated in the second row in Fig. S1. It should be noticed that $M_{x}^{2}(L)$ and $D^{2}(L)$ are calculated at different $q_{c}(L)$ as the crossing points are got from different dimensionless quantities. In the calculation of $M_{x}^{2}(L)$ we chose the corresponding $q_{c}(L)$ got from $U_{M}$ for two sizes $L$ and $2L$. As for the $D^{2}(L)$ the crossing points are located using $U_{D}$. After all the $M_{x}^{2}(L)$ and $D^{2}(L)$ are known for different sizes, the square of order parameters at critical points when $L\rightarrow\infty$ can be calculated with the second-order polynomial fitting. All the fitting parameters of $M_{x}^{2}(\infty)$ and $D^{2}(\infty)$ for different $\Delta$ are illustrated in Fig. 1 in the main text.
II Comparison of fitting quality

### v3

@@ III Disorder Operator
Once the transition points of each $\Delta$ are determined, we now carry out the analysis of the disorder operator upon them. As a non-local observable, disorder operator is defined as the expectation value of a symmetry transformation applied to a finite region in the statistical or quantum many-body systems of interest [49, 50, 51, 52, 53]. The design and implementation of the disorder operator and the analysis of its finite size scaling behavior have been successfully carried out in the situations of spontaneous symmetry breaking phase, quantum critical points, the symmetric phases with topological orders, symmetric mass generation transition and even the free fermion surface and interacting quantum critical Fermi surface systems [54, 55, 56, 57, 58, 11, 59, 15, 17, 60, 61].
@@
We have further computed a related quantity – the bipartite spin fluctuation [62]. At $\theta\to 0$ limit, the scaling of the disorder operator $\ln|X_{M}|$ has a similar behavior to that of the bipartite spin fluctuations, which is defined as
@@
Figure 4: Second Rényi entropy $S^{(2)}_{A}$ for $L\times L$ lattice with subsystem size of $L/2\times L$ deep in the AFXY phase of EPJQ model at $\Delta=0$ and q=0.3q=0.3.(a) $S^{(2)}_{A}$ versus system size $L$. The black curve is a fit to $S^{(2)}_{A}=aL+b\ln(L)+c$ and the fitting result is $S^{(2)}_{A}=0.1338(8)L+0.497(11)\ln(L)+0.0046(166)$. (b) $S^{(2)}_{A}-aL$ vs. $\ln(L)$ for different system sizes where we choose the fitted value $a=0.1338$. The black curve is a straight line of $y=0.497x+0.0046$.
@@ IV Entanglement Entropy
Next we investigate the EE at the transition points of various $\Delta$ of Eq. (1). To this end, we employ the nonequilibrium increment method [63, 64, 12, 65] within the framework of SSE-QMC simulation [43, 44] to determine the second Rényi entropy of the EPJQ model at phase transition points for various values of $\Delta$. Rényi entanglement entropy is defined as $S_{A}^{(n)}=\frac{1}{1-n}\ln\operatorname{Tr}\rho_{A}^{n}$ which can be reexpressed in the form of $S_{A}^{(n)}=\frac{1}{1-n}\ln\frac{Z_{A}^{(n)}}{Z^{(n)}}$ according to the replica trick [66]. The nonequilibrium method is based on Jarzynski’s equality [67], which relates the free energy difference between two systems with the total work done during a tuning process from one system to another. We regard the partition functions $Z_{A}^{(n)}$ and $Z^{(n)}$ as those of two different physical systems, then it is natural to apply the Jarzynski’s equality and design a tunning process between the two systems to calculate the Rényi entropies [63, 64]. In practice, we follow the incremental version [12, 65] of Ref. [64], and it can overcome the obstacles that the EE is in general an exponential observable [68, 69]. We conduct the simulation on a $L\times L$ square lattice with periodic boundary conditions, and fix with $\beta=L$ and choose the subregion A to be a $L\times L/2$ cylinder defined same as region $M$ in the right inset of Fig. 1 (a), which has no sharp corners on the entanglement boundary.
For $\Delta<1.0$, in the AFXY phase, the system spontaneously breaks the SO(2) spin symmetry and possesses one Goldstone mode. In this case, EE is expected to scale as [32]
@@
where $N_{g}$ represents the number of Goldstone modes. Eq. (9) has been extensively verified for both XY phase with one Goldstone mode and Néel ordered states with two Goldstone modes respectively [70, 71, 64, 65, 13, 72, 69, 73] . We first test our algorithm in the AFXY phase of the EPJQ model at $\Delta=0$ and $q=0.3$ deep in the AFMXY phase, with the results displayed in Fig. 4. We measure the second order Rényi EE for the $L\times L/2$ subsystem $A$ on a $L\times L$ square lattice with smooth boundary. The system sizes utilized are $L=8,12,16,\ldots,32$ and the temperature is fixed at $\beta=L$. As shown in Fig. 4(b), by fitting the finite-size EE data to Eq. (9), we observe a good agreement between our data with the scaling form of Eq. (9). Moreover, the log-coefficient obtained from curve fitting is $\frac{N_{G}}{2}=0.497\pm 0.011$, aligning well with the expectation of existence of one Goldstone mode, which corresponds to spontaneous SO(2) symmetry breaking. The analysis of the fitting quality can be found in the Appendix B.
@@
We then proceed to the phase transition points. Our extrapolated finite order parameter data in Fig. 1(b) shows that the first-order behavior is more pronounced as $\Delta$ decreases from 1 to 0. However, the relation between the scaling of entanglement entropy (EE) with the strengthen of first-order behavior remains unexplored. We measure the second Rényi EE at the transition points in Table. 1 for various values of $\Delta$ and system sizes $L=8,12,16,\ldots,44$ and fix $\beta=L$. The subsystem A is again chosen to be $L\times L/2$ cylinder withour corners on the boundary. As presented in Fig. 5, for all examined values of $\Delta$, we observe a clear logarithmic correction to the area law term in the scaling of EE (Eq. (9)). Additionally, as $\Delta$ decreases and the first-order behavior strengthens, the fitted log-coefficient $b$ increases. Interestingly, for $\Delta=0.1$ and $\Delta=0.3$, the fitted value of $b$ is even consistent with 0.5, which corresponds to the existence of a single Goldstone mode, the same as the scaling of EE deep inside the AFXY phase in Fig. 4. Our findings suggest that, for $\Delta\leq 0.3$, the logarithmic correction from single Goldstone mode can also be observed even in the mixed phase of Néel order and VBS state. For $\Delta>0.3$ and as $\Delta$ increases, the first-order behavior becomes weaker and weaker as the remaining order parameters suggest, it is very likely here the fitted $b$ suffers from strong finite-size effects and will converge to 0.5 as system size grows. Here with the limited system sizes, it is impossible to address this issue thoroughly. We leave this issue for further studies.
@@ V Discussion
Our work shows the two non-local measurements – the disorder operator and entanglement entropy – consistently exhibit anomalous scaling behaviors at the easy-plane deconfined quantum criticalities. By adjusting the anisotropy from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$), we find that as the first-order nature of the transition is amplified, the anomalous behavior for both disorder operator and entanglement entropy are also strengthened, and they resembles contributions from Goldstone modes. Interestingly, when $\Delta\leq 0.3$, the log coefficients in EE gradually converge to $b=0.5$ which is consistent with one Goldstone mode and SO(2) symmetry breaking at the first-order transition points. Our work thus provides strong evidence that the observed anomalous scaling behaviors of the entanglement measurements (disorder operator and EE) indeed come from the weakly-first-order nature of DQCPs of JQ model realizations both at $\Delta=1$ and $\Delta<1$.
Future research directions may include investigating the disorder operator and entanglement entropy scaling of JQn models with $n>3$, which have been shown to exhibit stronger first-order transitions [74, 39] as $n$ increases, similar to EPJQ models. Another avenue of interest is to explore the entanglement entropy scaling of an absolutely strong first-order transition, which involves a mixture of two phases with classical probabilities. Moreover, multipartite entanglement [75, 76, 77, 78] and higher-order Rényi entropies, entanglement negativity and entanglement spectrum [79, 80, 81] with QMC simulations at the DQCP also present intriguing topics for future investigation.
@@ Acknowledgement
We thank Meng Cheng, Cenke Xu, Anders Sandvik, Senthil Todadri, Subir Sachdev for valuable discussions on the related topic. NVM and ZYM acknowledge the earlier insightful discussions on the transitions of EPJQ model with Arnab Sen and Anders Sandvik. JRZ and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037- 22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A HKU703/22), the GD-NSF (No. 2022A1515011007) and the HKU Seed Funding for Strategic Interdisciplinary Research. Y.C.W. acknowledges the support from the Natural Science Foundation of China (Grant No. 12474216), the Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003), and the start-up funding and the High-Performance Computing Centre of Hangzhou International Innovation Institute of Beihang University. NVM acknowledges the National Natural Science Foundation of China (No. 12004020) and the Fundamental Research Funds for the Central Universities. We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.
Appendix A The criticality for different $\Delta$ in the EPJQ models
@@
Figure 6: (Upper panels) The crossing points $q_{c}(L)$ obtained from the curve of dimensionless Binder cumulants $U_{M}$(blue) and $U_{D}$(red) versus $q$ close to the phase transition points for two system sizes $L$ and $2L$. The red/blue curve is the fitting function with the scaling form of $q_{c}(L)$ in Eq. (5) excluding the crossings for small sizes with $\chi^{2}$ of the fitting becomes close to one. The data points presented at $1/L=0$ are the y-intecepts of fitting curves with the fitting errors which are given in Tab.1. (Lower panels) The value of order parameters for both AFXY and VBS phases at $q_{c}(L)$ obtained with corresponding Binder cumulants versus $1/L$ in the EPJQ models with different $\Delta$ for all computing sizes $L$. The red/blue curve is the second-order polynormial fitting function of all the data points with $\chi^{2}$ of the fitting close to one. The data points presented at $1/L=0$ are the y-intecepts of the fitting curves with errors illustrated in Fig. 1.
@@
The results in (b) of Fig. 1 in the main text present that the order parameters for both the AFXY and VBS phases converge to a non-zero value when $L$ goes to infinity for different $\Delta$ in the EPJQ models, which implies the possibility of first-order phase transitions in the EPJQ model. In this section we present the detail procedure in obtaining those converged order parameters at thermodynamic limit in Fig. 1(b). The binder cumulants $U_{M}$ and $U_{D}$ defined in Eq. (4) for the AFXY and VBS phases correspondingly are two dimensionless quantities that are commonly chosen in the FSS of calculating critical points and correlation length exponents. In this paper we study the critical behavior of the EPJQ model with different anisotropic value $\Delta$ with the help of $U_{M}$ and $U_{D}$ all the crossing points $q_{c}(L)$ of two simulated sizes $L$ and $2L$ got from $U_{M}$ and $U_{D}$ are shown in the first row in Fig. 6 for those different $\Delta$. Using the fitting form in Eq. (5) two $q_{c}(\infty)$ can be got, which are listed in Tab.1 in the main text. At each crossing point $q_{c}(L)$ we also calculate the square of order parameters defined in Eq. 2 and Eq. 3 as $M_{x}^{2}(L)$ and $D^{2}(L)$ illustated in the second row in Fig. 6. It should be noticed that $M_{x}^{2}(L)$ and $D^{2}(L)$ are calculated at different $q_{c}(L)$ as the crossing points are got from different dimensionless quantities. In the calculation of $M_{x}^{2}(L)$ we chose the corresponding $q_{c}(L)$ got from $U_{M}$ for two sizes $L$ and $2L$. As for the $D^{2}(L)$ the crossing points are located using $U_{D}$. After all the $M_{x}^{2}(L)$ and $D^{2}(L)$ are known for different sizes, the square of order parameters at critical points when $L\rightarrow\infty$ can be calculated with the second-order polynomial fitting. All the fitting parameters of $M_{x}^{2}(\infty)$ and $D^{2}(\infty)$ for different $\Delta$ are illustrated in Fig. 1 in the main text.
Appendix B Comparison of fitting quality